OpenStudy (selenagarcia889):

Use the Pythagorean Theorem to find the distance between (-3, 4) and (6, 1). Show all calculations.

OpenStudy (anonymous):

Do you know your equation?

OpenStudy (selenagarcia889):

Nope all it has is that and i dont understand it at all

OpenStudy (anonymous):

a^2+b^2=c^2 Would be your Equation for the Pythagorean Therom

OpenStudy (selenagarcia889):

Thanks

OpenStudy (shadowlegendx):

Do you know how to use it?

OpenStudy (selenagarcia889):

d=(x-x)+(y-y) RIght?

OpenStudy (shadowlegendx):

No, you can use the Pythagorean Theorem here

OpenStudy (shadowlegendx):

You just have to use a plot

OpenStudy (shadowlegendx):

Let (-3,4) be A and let (6,1) be B

OpenStudy (shadowlegendx):

Graph the two points

OpenStudy (shadowlegendx):

Let me know if you need help